Lisaneedham:

What is the inverse of the function f(x) = 4x + 8? A. h(x) = one-quarterx – 2 B. h(x) = one-quarterx + 2 C. h(x) = one-halfx – 2 D. h(x) = one-halfx + 2

supie:

Sorry it said i was typing but i was afk but My math is..... The function: f(x) = 4x+8 rewrite the function with y=f(x) y=4x+8 Now we have to solve for x so, y-8 = 4x 2= y-8/4 x=y/4-2 Now change the variables y=x/4 Last we rewrite the function h(x) = y h(x) = x/4 -2 So, the inverse function is h(x) = x/4-2 That should be right.

Lisaneedham:

ok thanks

supie:

Np.